# Jose Tupayachi

### v1

@@ 2.3 Our Contribution
Recognizing the existing research gap, our paper proposes a TW-GCN, which consists of a GCN model to learn the spatial features of station-level geographic characteristics, and a time series model. This is wrapped by temporal component carrying historical datasets including weather, and traffic. Subsequently, a construct of a stacked predictive model integrates these insights to forecast the EV charging demand. Our contributions are threefold as follows:
  • We implement real-world traffic and weather features on temporal-based GCN method and analyze the impact on the forecasting behavior while benchmark against baseline methods.
  • We leverage proprietary EV transaction data provided by a leading charging network operator in Tennessee to examine charging behavior in a region that lags in electric vehicle adoption.
  • We conduct computational experiments that integrate time series models with graph convolutional networks to enhance the accuracy of EV charging demand predictions.
@@ 3 Datasets
In order to effectively model and predict charging station energy consumption, we compile and integrate diverse datasets capturing temporal, spatial, and contextual factors. This study forecasts energy consumption (kWh) at charging stations by leveraging various influencing variables (see Table 2). Temporal attributes Zhang et al. (2025), such as start_date_time, reveal usage patterns (Senol et al., 2023; Dominguez-Jimenez et al., 2020), while geographic coordinates (station_latitude, station_longitude) capture spatial relationships (Li et al., 2018). Additionally, traffic metrics, weather conditions, and POI data are incorporated to enrich the modeling context. These temporal, spatial, and contextual attributes are grounded in real-world charging behavior, motivating the use of the ChargePoint dataset to analyze station usage and infrastructure patterns.
@@ 3.1 ChargePoint Dataset & Stations
EV adoption has accelerated in recent years, fueling the expansion of charging infrastructure to meet growing demand. Reliable and comprehensive datasets on station usage are crucial for understanding user behavior, optimizing network deployment, and enhancing operational efficiency. The ChargePoint network, the largest independently owned EV charging provider in North America and Europe, operating in more than 14 countries, offers valuable insights into charging patterns and infrastructure utilization (ChargePoint Investors, 2024). As of January 2025, it accounts for 61% of publicly available AC charging ports in North America and provides access to over one million charging locations when including roaming partners (U.S. Securities and Exchange Commission, 2025; ChargePoint Investors, 2024). In Tennessee, 86 ChargePoint stations represent about 7.5% of the state’s 1,142 public charging locations (U.S. Department of Energy, 2025).
The dataset used in this study comprises detailed session-level logs from these 86 stations over a 13-month period. Each log captures key variables such as charging duration, energy dispensed, and station-specific metadata. To support our modeling framework, the dataset is further enhanced with the weather, traffic, and points of interest features, enabling a more nuanced analysis of spatial and temporal demand patterns. As a sample, Figure 1 illustrates the aggregated energy consumption of stations grouped by zip code. Building on this spatial aggregation, the electric vehicle charging behavior also requires examining temporal usage patterns, which adds further complexity to the analysis.
Figure 1: Energy consumption in kWh for the year 2024-2025. Colored regions represent the zip codes with the highest energy usage.
@@ 3.2 Points of Interest
Understanding the factors that shape charging demand requires considering the built environment around stations. In particular, nearby POIs can significantly influence whether certain stations are used more frequently than others (Pagany et al., 2019). Using a minimum-distance criterion (Straka et al., 2020), we incorporate POIs into various variables, including spatial patterns derived from POI proximity. This integration helps to better capture spatial heterogeneity in charging demand. We selected locations within a 500-meter radius of each station to capture local demand patterns while maintaining spatial specificity (see Table 2 in Appendix). A 500-meter buffer, commonly used in urban studies, captures the typical walkable distance for EV users seeking nearby amenities, enabling analysis of spatial factors influencing charging demand Ren and Sun (2025). While POI information reflects potential interactions with demand estimates, factors such as off-street parking availability may respond to local conditions like weather Lyu et al. (2024). In our study, we collected POI information from OpenStreetMap (OpenStreetMap, 2025).
@@ 3.3 Weather and Traffic Information
Environmental factors such as weather and traffic conditions significantly influence EV charging behavior and station performance. In our work, we selected weather features including temperature, pressure, and humidity, as these directly affect the performance of the lithium-ion batteries used in EVs. For example, batteries experience reduced efficiency and accelerated degradation in cold environments (Wang et al., 2025; Senol et al., 2023), which can increase charging duration and, in turn, station usage. Conversely, hot climates, particularly in arid regions, create additional challenges for battery efficiency and station operations (El Hafdaoui et al., 2023).
In addition, we considered traffic indicators such as average speed and congestion within the EV corridor. These factors capture the effect of bottleneck-constrained roads, where structural and dynamic interactions can significantly alter charging behavior (Wang et al., 2021; Huang et al., 2016).
Weather-related variables such as temperature, humidity, and precipitation were obtained from Weather Underground (Underground, 2025). Transportation-related variables were derived from Traffic Message Channels (TMCs), which encode real-time traffic events including congestion, accidents, and road closures using standardized location and event codes (International Organization for Standardization, 2013; Wright and Rehborn, 2008). These features enable dynamic rerouting and optimized traffic flow within EV corridors by incorporating up-to-date traffic patterns. The traffic data was sourced from the Regional Integrated Transportation Information System (RITIS) (University of Maryland, 2025).
@@ 4 Proposed Methodology
Several methods explore how to address the complex demand prediction of electric vehicles. Authors, have coupled weather, traffic, and events data to predict EV session duration and energy consumption using traditional machine learning methods (Shahriar et al., 2021). In the following years, methods have added temporal capabilities, leading to improvements in the robustness of the conditions for EV charging demand (Chen et al., 2023). Most methods do not consider spatial relationships while including temporal capabilities paired with traffic, weather and point of interest characteristics (Chen et al., 2023; Hüttel et al., 2021) and Table 1 provides a description of similar approaches aimed at analyzing this information.
To capture the complex relationships between entities such as charging stations connected by traffic patterns or spatial proximity, graph-based learning provides a natural framework. In this paradigm, each node $v\in\mathcal{V}$ in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{F}$, where $F$ denotes the number of features. Here, $\mathcal{V}$ represents the set of nodes (e.g., charging stations), and $\mathcal{E}$ represents the set of edges that capture relationships or interactions (e.g., spatial proximity or traffic influence). These vectors capture key attributes of the entities represented by the nodes, enabling the model to incorporate both structural and contextual information. In our study, node features serve as the primary input to the Network and capture intrinsic attributes of charging stations, including geographic location, capacity, historical energy usage, and operational metrics like session duration and charger availability. These vectors enable the model to learn contextualized representations through iterative message passing and neighbor aggregation. Node features vary by domain for instance, voltage and power measurements in power grids (Suri and Mangal, 2025), or socio-demographic, land use, and traffic data in EV infrastructure modeling (Batic et al., 2025). Message passing layers iteratively update node features by incorporating information from adjacent nodes, allowing each station to reflect its local network context.
[…]
@@
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$, individual vertex in the set of nodes $\mathcal{V}$ of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and $f_{\theta}^{(\ell+1)}$ is a learnable function that aggregates the node features with those of its neighbors. We adopt the GCN formulation introduced by Kipf and Welling (2016), in which the layer-wise propagation rule (Equation 2) is defined as:
@@ 4.2 Temporal Stacking
where $N$ is the number of nodes (stations), $H$ is the hidden feature dimension, and $T$ is the number of time steps. In order to introduce the temporal components, the GCN outputs are passed to a time series model (e.g., LSTM). Represented by: $\mathbf{Z}^{\text{geo}}\in\mathbb{R}^{T\times N\times H},\quad\mathbf{Z}^{\text{dem}}\in\mathbb{R}^{T\times N\times H}$
@@
  • Geographical features ($\mathbf{Z}^{\text{geo}}$): variables that are tied to location and remain relatively static, such as POIs.
@@
where the index $j$ denotes the cluster or region identified by K-means, and $\mathbf{C}_{j}$ is the centroid of the $j$-th cluster. For each node $i$, this expression assigns it to the region $j$ whose centroid $\mathbf{C}_{j}$ is closest to the node’s feature vector $\mathbf{Z}_{T}(i)$ at the last time step $T$, effectively grouping nodes into regions based on feature similarity. A MLP is then applied to model the nonlinear relationship between the shared feature representation $\mathbf{Z}$ and the energy consumption $\hat{\mathbf{Y}}_{i}$ for each region, enabling cluster-specific predictive behavior. MLPs serve as fundamental components of state-of-the-art transformer architectures, where they operate alongside attention mechanisms to capture complex dependencies over learned representations. The final output applies a ReLU activation to ensure non-negative kWh values, aligning with physical constraints (Equation 7).
@@ 4.3 Graph Generation
In order to effectively capture both spatial and temporal dependencies within the data. We construct spatial and temporal relationships between nodes, using two distinct adjacency matrices: One for geographic proximity and another for dynamic time warping-based similarity in demand patterns. The geospatial adjacency matrix is defined as follows: given a set of $N$ nodes, each associated with spatial coordinates $(\phi_{i},\lambda_{i})$ (latitude and longitude), we compute the pairwise distance using the Haversine formula (Algorithm 2). The geospatial adjacency matrix $A^{\text{geo}}\in\mathbb{R}^{N\times N}$ is then constructed by applying a threshold $\tau$ or a Gaussian kernel to these distances, ensuring that physically proximate nodes are more strongly connected (see Figure 2 for the architecture overview). Likewise, the temporal similarity matrix is quantify similarities in demand patterns using the DTW distance, as defined in Algorithm 1, between historical demand time series for each node pair (Equation 8):
@@ 4.4 Clustering
We employ the K-means clustering algorithm on station-level data to uncover hidden spatial patterns in charging behavior. Prior research demonstrates that clustering effectively identifies distinct user groups with similar behavioral patterns Nespoli et al. (2023); Genov et al. (2024), particularly in energy charging habits (Al-Ogaili et al., 2019). These clusters often reflect underlying geographic context, enabling more accurate forecasting and tailored strategies.
@@ 5 Computational Experiments Setup
We conduct an ablation study by varying: lag hours with values 1, 3, and 6. This parameter delimits how many hours in advance our prediction will forecast, a number of geographic clusters (NUMCLUSTERSLIST with 5 and 10), models (MODELSLIST with ‘LSTM’, ‘GRU’, ‘1DCNN’, ‘RNN’) representing what models are used, sequence lengths to create the necessary time windows to (SEQLENGTHLIST with 4, 8, and 12), and alpha values (ALPHALIST with 0.33, 0.5, and 0.66) weight probability during the concatenation process. Boolean parameters include traffic, weather, and adjacency features (TRAFFICFEATURE, WEATHERFEATURE, ADJACENCY set to True), and the number of epochs (NUMEPOCHS set to 125). The method was initially tested following the approach of Wang et al. (Wang et al., 2023b, a) and later extended to publicly available datasets, including the Boulder, Colorado EV charging dataset [^1] https://open-data.bouldercolorado.gov/maps/95992b3938be4622b07f0b05eba95d4c_0/explore. Subsequently, we utilized a proprietary ChargePoint dataset centered in the region of study. The Boulder dataset provides spatiotemporal records of charging events, including start and end times, location, energy delivered, and environmental impact metrics, enabling the modeling of temporal and spatial dependencies between stations.

### v3

@@ 2.3 Our Contribution
Recognizing an existing research gap, this paper proposes a TW-GCN, which combines a GCN model for capturing spatial features of station-level geographic characteristics with a temporal model that processes historical time series data, including weather and traffic information. This stacked architecture integrates spatial and temporal predictors to forecast EV charging demand. Our contributions are threefold:
  • We incorporate real-world traffic and weather features into a temporal GCN framework and analyze the impact on forecasting performance, benchmarking against baseline methods.
  • We leverage proprietary EV transaction data from a leading charging network operator in Tennessee to study charging behavior in a region characterized by slower yet steadily growing EV adoption.
  • We perform computational experiments that combine time series models with graph convolutional networks to improve the accuracy of EV charging demand predictions.
@@ 3 Datasets
To effectively model and predict charging station energy consumption, we compile and integrate diverse datasets capturing temporal, spatial, and contextual factors. We forecast energy consumption (kWh) at charging stations by leveraging these various influencing variables illustrated in Table 2. Temporal attributes (zhang2025high), such as start_date_time, reveal usage patterns (senol2023electric; dominguez2020seasonality), while geographic coordinates (station_latitude, station_longitude) capture spatial relationships (li2018planning). Additionally, traffic-related metrics, weather conditions, and POI data are incorporated to provide a richer modeling context. These temporal, spatial, and contextual attributes reflect real-world charging behavior, motivating the use of a commercial dataset to analyze station usage and infrastructure patterns.
Table 2: Features and Definitions per category.
@@ 3.1 ChargePoint Dataset & Stations
EV adoption has accelerated in recent years, fueling the expansion of charging infrastructure to meet growing demand. Reliable and comprehensive datasets on station usage are crucial for understanding user behavior, optimizing network deployment, and enhancing operational efficiency. The ChargePoint network, the largest independently owned EV charging provider in North America and Europe, operating in more than 14 countries, offers valuable insights into charging patterns and infrastructure utilization (investors2024globalreach). As of January 2025, it accounts for 61% of publicly available AC charging ports in North America and provides access to over one million charging locations when including roaming partners (sec2025filing). In Tennessee, 86 ChargePoint stations represent about 7.5% of the state’s 1,142 public charging locations (doe2025afdc).
The dataset used in this study comprises detailed session-level logs from these stations over 13 months. Each log captures key variables such as charging duration, energy dispensed, and station-specific metadata. To support our modeling framework, the dataset is further enhanced with the weather, traffic, and points of interest (POI) features, enabling a more nuanced analysis of spatial and temporal demand patterns.
@@ 3.2 Points of Interest
Understanding the factors that shape charging demand requires considering the built environment around stations. In particular, nearby POIs can significantly influence whether certain stations are used more frequently than others (pagany2019electric). Using a minimum-distance criterion (straka2020predicting), we incorporate POIs into various variables, including spatial patterns derived from POI proximity. This integration helps to capture spatial heterogeneity in charging demand. We selected locations within a 500-meter radius of each station to capture local demand patterns while maintaining spatial specificity, as shown in Table 2. A 500-meter buffer, commonly used in urban studies, captures the typical walkable distance for EV users seeking nearby amenities, enabling analysis of spatial factors influencing charging demand (ren2025predicting). While POI information reflects potential interactions with demand estimates, factors such as off-street parking availability may respond to local conditions like weather (lyu2024short). In our study, we collected POI information from OpenStreetMap (OpenStreetMap).
@@ 3.3 Weather and Traffic Information
Environmental factors such as weather and traffic conditions significantly influence EV charging behavior and station performance. In our work, we selected weather features including temperature, pressure, and humidity, as these directly affect the performance of the lithium-ion batteries used in EVs. For example, batteries experience reduced efficiency and accelerated degradation in cold environments (wang2025electric), which can increase charging duration and, in turn, station usage. Conversely, hot climates, particularly in arid regions, create additional challenges for battery efficiency and station operations (el2023impact).
In addition, we considered traffic indicators such as average speed and congestion within the EV corridor. These factors capture the effect of bottleneck-constrained roads, where structural and dynamic interactions can significantly alter charging behavior (wang2021electric).
Weather-related variables such as temperature, humidity, and precipitation were obtained from Weather Underground (WeatherUnderground). Transportation-related variables were derived from Traffic Message Channels (TMCs), which encode real-time traffic events including congestion, accidents, and road closures using standardized location and event codes (iso14819; wright2008tmc). These features enable dynamic rerouting and optimized traffic flow within EV corridors by incorporating up-to-date traffic patterns. The traffic data was sourced from the Regional Integrated Transportation Information System (RITIS) (ritis).
@@ 4 Proposed Methodology
Studies (shahriar2021prediction; huttel2021forecasting; chen2023hybrid) have examined methods to address the complex task of forecasting EV charging demand. These approaches commonly integrate contextual factors such as weather, traffic, and event data to predict session duration and energy consumption using traditional machine learning models. While incorporating temporal dependencies improves predictive robustness, most methods do not account for spatial relationships, even when temporal features are combined with traffic, weather, and POI information.
To capture the complex relationships between entities such as charging stations connected by traffic patterns or spatial proximity, graph-based learning provides a natural framework. In this paradigm, each node $v\in\mathcal{V}$ in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{F}$, where $F$ denotes the number of features. Here, $\mathcal{V}$ represents the set of nodes (e.g., charging stations), and $\mathcal{E}$ represents the set of edges that capture relationships or interactions (e.g., spatial proximity or traffic influence). These vectors capture key attributes of the entities represented by the nodes, enabling the model to incorporate both structural and contextual information.
In our study, node features serve as the primary input to the Network and capture intrinsic attributes of charging stations, including geographic location, capacity, historical energy usage, and operational metrics like session duration and charger availability. These vectors enable the model to learn contextualized representations through iterative message passing and neighbor aggregation. Node features vary by domain, for instance, voltage and power measurements in power grids (suri2025gnn), or socio-demographic, land use, and traffic data in EV infrastructure modeling (batic2025evgnn). Message passing layers iteratively update node features by incorporating information from adjacent nodes, allowing each station to reflect its local network context.
@@
Where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$, individual vertex in the set of nodes $\mathcal{V}$ of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and $f_{\theta}^{(\ell+1)}$ is a learnable function that aggregates the node features with those of its neighbors. We adopt the GCN formulation introduced by kipf2016semi, in which the layer-wise propagation rule (Equation 2) is defined as:
@@ 4.2 Temporal Stacking
where $N$ is the number of nodes (stations), $H$ is the hidden feature dimension, and $T$ is the number of time steps. To introduce the temporal components, the GCN outputs are passed to a time series model (e.g., LSTM). Represented by: $\mathbf{Z}^{\text{geo}}\in\mathbb{R}^{T\times N\times H},\quad\mathbf{Z}^{\text{dem}}\in\mathbb{R}^{T\times N\times H}$
@@
  • Geographical features ($\mathbf{Z}^{\text{geo}}$): variables that are tied to a location and remain relatively static, such as POIs.
@@
where the index $j$ denotes the cluster or region identified by K-means, and $\mathbf{C}_{j}$ is the centroid of the $j$-th cluster. For each node $i$, this expression assigns it to the region $j$ whose centroid $\mathbf{C}_{j}$ is closest to the node’s feature vector $\mathbf{Z}_{T}(i)$ at the last time step $T$, effectively grouping nodes into regions based on feature similarity. An MLP is then applied to model the nonlinear relationship between the shared feature representation $\mathbf{Z}$ and the energy consumption $\hat{\mathbf{Y}}_{i}$ for each region, enabling cluster-specific predictive behavior. MLPs serve as fundamental components of state-of-the-art transformer architectures, where they operate alongside attention mechanisms to capture complex dependencies over learned representations. The final output applies a ReLU activation to ensure non-negative kWh values, aligning with physical constraints (Equation 7).
@@ 4.3 Graph Generation
To effectively capture both spatial and temporal dependencies within the data, we construct relationships between nodes using two distinct adjacency matrices: one representing geographic proximity and another capturing the DTW similarity in demand. The geospatial adjacency matrix is defined as follows: given a set of $N$ nodes, each associated with spatial coordinates $(\phi_{i},\lambda_{i})$ (latitude and longitude), we compute pairwise distances using the Haversine formula (Algorithm 2). The resulting geospatial adjacency matrix $A^{\text{geo}}\in\mathbb{R}^{N\times N}$ is constructed by applying a threshold $\tau$ to ensure geographically proximate nodes are strongly connected (see Figure 1 for an overview of the architecture). Similarly, the temporal similarity matrix quantifies the similarity of demand patterns by computing the DTW distance, as defined in Algorithm 1, between the historical demand time series of each pair of nodes (Equation 8).
[…]
Algorithm 1 Dynamic Time Warping (DTW)
1:Two sequences $s=[s_{1},s_{2},\ldots,s_{n}]$, $t=[t_{1},t_{2},\ldots,t_{m}]$
2:DTW_dist – the DTW distance between $s$ and $t$
3:Initialize array DTW[0..n, 0..m] with $\infty$
4:DTW[0, 0] $\leftarrow 0$
5:for $i\leftarrow 1$ to $n$ do
6:  for $j\leftarrow 1$ to $m$ do
7:   $cost\leftarrow d(s_{i},t_{j})$
8:   DTW[$i,j$] $\leftarrow cost+\min(\texttt{DTW}[i-1,j],\texttt{DTW}[i,j-1],\texttt{DTW}[i-1,j-1])$
9:  end for
10:end for
11:return DTW[n, m] $\triangleright$ Final DTW distance
Algorithm 2 Haversine Distance
1:lat1, lon1, lat2, lon2 – coordinates in degrees
2:d – Haversine distance in kilometers
3:Convert lat1, lon1, lat2, lon2 from degrees to radians
[…]
9:d $\leftarrow R\cdot c$
10:return d
@@ 4.4 Clustering
We employ the K-means clustering algorithm on station-level data to uncover hidden spatial patterns in charging behavior. Prior research demonstrates that clustering effectively identifies distinct user groups with similar behavioral patterns (nespoli2023user; genov2024forecasting), particularly in energy charging habits (al2019review). These clusters often reflect underlying geographic context, enabling more accurate forecasting and tailored strategies.
@@ 5 Computational Experiments
We employ several deep learning architectures to model temporal dependencies and nonlinear dynamics in EV charging demand. The RNN serves as a foundational model capable of capturing sequential dependencies through recurrent connections. Building upon this, the LSTM network introduces gating mechanisms to mitigate vanishing gradient issues and effectively learn long-term temporal relationships. Similarly, the GRUs simplify the LSTM structure while maintaining comparable performance by combining the forget and input gates. One-Dimensional Convolutional Neural Networks (1DCNNs) leverage convolutional filters to extract local temporal patterns from time-series data, offering computational efficiency and robust feature extraction. Together, these architectures form the core of our modeling framework, enabling a comprehensive exploration of temporal dynamics in EV charging demand. Finetuning parameter grid is found and optimization parameter are found under Table 3.
Table 3: Summary of TW-GCN model parameters, input dimensions, and parameter statistics under optimal configuration.
Hyperparameter Type Default Description
node_features int 18 Number of dynamic input features per node.
hidden_size int 85 Dimensionality of hidden representations for GCN and temporal layers.
num_regions int 85 Number of distinct geographic or functional regions.
alpha float 0.1 Fusion weight between geographic and demand graph outputs: $x=\alpha x_{\text{geo}}+(1-\alpha)x_{\text{dem}}$.
num_clusters int 10 Number of K-Means clusters for regional grouping.
model_type str ’LSTM’ Temporal model type. Options: {’GRU’, ’LSTM’, ’BiLSTM’, ’RNN’, ’1DCNN’}.
gcn_geo, gcn_dem GCNConv Graph convolution layers for geographic and demand graphs. Input: $(F)$, Output: hidden_size.
temporal_model_geo, temporal_model_dem $n$.Module LSTM Recurrent models for temporal embedding in geographic and demand subgraphs.
region_mlp $n$.ModuleList Linear() Region-specific output heads for local predictions.
relu Activation ReLU() Non-linearity ensuring non-negative outputs.
n_init int 10 Number of random initializations for K-Means clustering.
batch_first bool True Treats input tensors as (batch, time, features).
kernel_size int 3 1D convolution kernel size (1DCNN temporal model).
padding int 1 Padding size to preserve sequence length in 1DCNN mode.
Model Dimensions and Input Settings per Fold
T (time steps) int 7021 Total number of temporal steps.
N (locations) int 85 Number of nodes or spatial regions.
F (features) int 18 Number of dynamic input features per node.
GCN output features int 18 Output dimensionality of graph convolutional embeddings.
hidden_state shape Tensor [batch, 85, 85] Intermediate representation for temporal modeling.
batch size int 32 Number of samples processed per training batch.
epochs int 100 Total number of training epochs | Subject to early stoping.
learning rate float 0.001 Step size used by the optimizer for gradient updates.
Input and Batch Tensor Shapes
adj_geo batch Tensor [32, 4, 85, 85] Geographic adjacency matrix batch shape.
adj_dem batch Tensor [32, 4, 85, 85] Demand adjacency matrix batch shape.
feature batch Tensor [32, 4, 85, 18] Node-level feature input tensor.
target batch Tensor [32, 4, 85, 1] Ground truth target tensor.
Layer and Parameter Summary
Total parameters int 52,730 Includes GCN, recurrent, and MLP layers.
Trainable params int 52,730 All parameters are trainable.
Input size MB 8.64 Memory footprint of model inputs.
Forward/backward size MB 31.22 Memory usage during forward and backward propagation.
Parameter size MB 0.21 Total memory used by model parameters.
Estimated total size MB 40.07 Approximate runtime memory usage per batch.
@@
We conduct an ablation study by varying lag hours with values 1, 3, and 6. This parameter delimits how many hours in advance our prediction will forecast, several geographic clusters (NUMCLUSTERSLIST with 5 and 10), models (MODELSLIST with ‘LSTM’, ‘GRU’, ‘1DCNN’, ‘RNN’) representing what models are used, sequence lengths to create the necessary time windows to (SEQLENGTHLIST with 4, 8, and 12), and alpha values (ALPHALIST with 0.33, 0.5, and 0.66) weight probability during the concatenation process. Boolean parameters include traffic, weather, and adjacency features (TRAFFICFEATURE, WEATHERFEATURE, ADJACENCY set to True), and the number of epochs (NUMEPOCHS set to 125). The stacked architecture (wang2023short) was developed following the methodological framework proposed by wang2023predicting. The design of the network layers and the tuning of hyperparameters were informed by these prior works to maintain methodological consistency and ensure optimal model performance. Building on this foundation, the architecture was further adapted and applied to a proprietary ChargePoint dataset representing the study region.